### v5

@@ 3 Algorithm and Analysis
3.1 Algorithm Components
Our method uses $\hat{g}(\boldsymbol{X}_{kl})$ to prioritise cluster-layer proposals to update (i.e. pop), and $G(\boldsymbol{X}_{l})$ to facilitate local optimisation by identifying and cutting off further exploration of dominated solutions. It iterates from layer 00 (i.e. leaf nodes) through to $l_{\text{max}}$, storing unique and non-dominated cluster mappings on separate search branches. The algorithm terminates if, given current cluster-layer proposals on the queue and additive costs, the current costs of these potential solutions exceed the least-cost solutions found so far. Potentially, near-optimal solutions can be returned at earlier iterations of the search.
3.1.1 Data Structures and Initialisations
@@
Table 2: Definition of data structures. We use \$currency-dollar\\$\$ to denote a named field within a data structure; for example, the layer field $l$ within a popped queue entry $q$ is denoted $q\$l$.
@@
$\text{Br}_{\text{lk}}\in\\ \mathbb{R}^{B\times l_{\text{max}}\times B\times l_{\text{max}}}\end{aligned}$ $\text{Br}_{\text{lk}}[i,l,j,l]=1$ indicates a link at layer $l$ between branches $i$ and $j$.
@@
$Q$, $Q^{\prime}$ A queue with entries $q=\{b,u(\boldsymbol{X}_{kl}),l,\hat{g}(\boldsymbol{X}_{kl}),G(\boldsymbol{X}_{l})\}$. We denote $Q^{\prime}$ as eligible entries for updating (popping) $q\in Q:\text{Br}_{\text{pl}}[b]\geq q\$l$, aka activated branches.
@@
3.1.2 Pop Queue or Activate Branch
Potentially, $Q$ could be treated as a heap to maximise computational efficiency [16]. We use $Pop()$ to refer to removing the lowest cost queue entry, choosing between equicost entries arbitrarily (Component Pop Queue or Activate Branch). If there are no eligible entries on the queue, i.e. no branch has $\text{Br}_{\text{pl}}\geq l$, then a branch is chosen for activation. This makes queue entries on that branch potentially eligible. Activation randomly switches between using $Peek()$ to return, but not remove, the lowest cost entry, or $Peek^{\prime}()$ to return an entry at the lowest layer on the queue, choosing arbitrarily between entries at the same layer. This facilitates switching between best-first and breadth-first search (Theorem 3.1). After activation, DCMAP tries again to pop the queue if it is not empty; the algorithm only proceeds to the next step (Component Prune Linked Dominated Branches) if a queue entry was successfully popped.
@@ 3.1.2 Pop Queue or Activate Branch
5:        $\text{Br}_{pl}(q\$b)\leftarrow|\text{Br}_{pl}(q\$b)|$
3.1.3 Prune Linked Dominated Branches
@@
3.1.4 Propose Cluster Mappings
The requirement for contiguous clusters (Definition 2.3) means that a node can only be mapped to a new cluster $u_{\text{new}}$, or the cluster of one or more child nodes $\bf{u}_{\text{exist}}$. For the former, we can ensure uniqueness and avoid accidentally mapping to an existing cluster by associating a unique label $1..n$ with each node. For the latter, define a priori a $n\times n$ matrix $S:S[i,j]=\{0,1\}$, where $1$ indicates that node $X_{i}$ can potentially be mapped to the same cluster as node $X_{j}$. To ensure contiguity, Algorithm 2 works backward by identifying, for a given node $X_{i}$, which nodes $X_{j}$ at layer $l(X_{i})$ or earlier layers can be in the same cluster as $X_{i}$ that are connected by a common parent node. Note that descendants are child nodes, child of child nodes and so on. For node $X_{j}$, $u(X_{i^{\prime}}):S[i^{\prime},j]=1$ are all the node clusters that $X_{j}$ could be mapped to. To preserve contiguity, node $X_{j}$ must be such that there are no paths $\boldsymbol{X}_{p}$ following the directed arcs between $X_{i}$ and $X_{j}$ containing any other nodes i.e. $\boldsymbol{X}_{p,ij}=\{X_{i},X_{j}\}$ or $\boldsymbol{X}_{p,ij}=\varnothing$. The former implies $X_{j}$ is a parent of $X_{i}$, and the latter is where both nodes are at the same layer, $l(X_{j})=l(X_{i})$.
@@ 3.1.4 Propose Cluster Mappings
When applied iteratively and incrementally from layer 00 to $l_{\text{max}}$ (Fig. 2), all feasible contiguous cluster mappings (Definition 2.3) can be generated using matrix $S$ (Algorithm 2) to identify all nodes $X_{j}$ that can be mapped to the same cluster as node $X_{i}$.
@@
The assumption of iterating incrementally over $l=0,1,\ldots,l_{\text{max}}$ means that at layer $l$, only nodes at layer $l$ or earlier need to be considered for mapping to the same cluster. Additionally, Algorithm 2 only allows nodes $X_{i}$ and $X_{j}$ to be mapped to the same cluster if they are directly connected (i.e. a parent or child of the parent) with no paths going through an intermediary node; this guarantees contiguity, even if $X_{i}$ and $X_{j}$ are not in successive layers (Remark 2.1). At each iteration, feasible cluster mappings for node $X_{j}$ are the clusters of nodes $X_{i}$ associated with non-zero rows of $S[,j]$, which might be in the same layer or at a child layer. By Definition 2.1 of layers, there must exist at least one parent-child path between each node in one layer and the previous layer, hence, there are no contiguous cluster mappings that cannot be generated by iteratively and incrementally applying $S$ from $l=0$ to $l=l_{\text{max}}$. ∎
@@
$\footnotesize\begin{bmatrix}&A&B&C&D&E&F&G\\ A&0&0&0&0&0&0&0\\ B&0&0&0&0&0&0&0\\ C&0&0&0&0&0&0&0\\ D&1&1&0&0&0&0&0\\ E&0&0&1&0&0&0&0\\ F&1&0&0&0&0&0&1\\ G&0&0&0&1&1&0&0\\ \end{bmatrix}$ (13)
For instance, node $F$ (row $F$) could share the same cluster as nodes $A$, $D$ and itself (Algorithm 2). Thus, when updating (popping) node $D$ (column $D$) as the search progresses from layer 00 to $2$, it could be mapped to the current cluster of node $F$ or $G$ or itself. Locally, this could lead to contiguous clusters involving D: $(A,F,D,G)$, $(A,F,D)$, $(F,D)$, $(D,G)$, $(D)$.
The above constrained space of feasible cluster mappings could be pre-computed at initialisation and are enacted in Component Update Possible Cluster Mappings and Queue through $\hat{\textrm{U}}$. Component Propose Cluster Mappings generates from $\hat{\textrm{U}}$ and particular $k$ and $l$ the cluster-mappings for cost evaluation at each iteration. Intuitively, for a given branch and layer $l$, $\boldsymbol{Z}_{1}$ represents the nodes that cannot be in any other cluster $k^{\prime}\neq k$, $\boldsymbol{Z}_{3}$ are nodes that have already been mapped to another cluster $k^{\prime}$, and $\boldsymbol{Z}_{2}$ are the nodes that could be in $k$ or another cluster $k^{\prime}$. Each of the $N$ unique cluster mapping combinations are stored on separate branches, producing $N-1$ new branches, done by updating data structures and duplicating queue entries from the current branch $b$ for new branches $b^{\prime}$.
@@
3:    $\left\{\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{N}\right\}\leftarrow\textsc{generatecombos}(\boldsymbol{Z},\text{U},\hat{\text{U}})$
@@
6:        Duplicate $q^{\prime}\in Q:q^{\prime}\$b=b,q^{\prime}\$l=l$ with $q^{\prime}\$b=b^{\prime}$
7:procedure generatecombos($\boldsymbol{Z},\text{U},\hat{\text{U}}$)
8:    $\boldsymbol{Z}_{2}\leftarrow\boldsymbol{Z}\setminus\{\boldsymbol{Z}_{1},\boldsymbol{Z}_{3}\}:\bigl{(}\text{U}[b,\boldsymbol{Z}_{3}]\neq u(\boldsymbol{X}_{kl}),U[b,\boldsymbol{Z}_{3}]\neq 0\bigr{)},\bigl{(}\forall\hat{\text{U}}[b,u\neq u(\boldsymbol{X}_{kl}),\boldsymbol{Z}_{1}]=0\bigr{)}$
9:    $\boldsymbol{X}_{i=1..N}\leftarrow\Bigl{\{}\boldsymbol{z}_{j}\cup\boldsymbol{Z}_{1}\text{ }\forall\boldsymbol{z}_{j}\in\text{Combinations}{\boldsymbol{Z}_{2}\choose 1..|\boldsymbol{Z}_{2}|},\boldsymbol{Z}_{1}\Bigr{\}}$
@@
$\hat{\textrm{U}}\left[b,l(\boldsymbol{X}_{i})=l\right]=\begin{bmatrix}k&X_{1}&X_{2}&X_{3}\\ 2&0&1&1\\ 3&1&1&1\\ \end{bmatrix}$ (14)
@@ 3.1.6 Layer Updates
3.1.7 Update Possible Cluster Mappings and Queue
@@
3.2 Theoretical Analysis
This section proves analytically that DCMAP is optimal in the sense that it finds all least cost cluster mappings, and does so such that an equally informed algorithm cannot do so in fewer iterations. This is achieved through conditional optimality (derived in Section 3.2.1), the exploration of all possible contiguous cluster mappings through branching (Section 3.2.2), and combined to show the optimality of DCMAP through evaluating non-dominated branches, i.e. potential cluster mapping solutions, at the earliest opportunity (Section 3.2.3). An additional property is derived for super additivity of cost functions for a heuristic to further speed up the search (Supplementary Information 6).
3.2.1 Conditional Optimality
@@
Following Definition 2.6, we say that $U^{*}_{<l}$ dominates $U^{i}_{<l}$ if $G^{*}(\boldsymbol{X}_{l}|U_{l},U^{*}_{<l})<G(\boldsymbol{X}_{l}|U_{l},U^{i}_{<l})$.
@@
${\sum_{k=1}^{m_{l+1}}{c\Bigl{(}\boldsymbol{X}_{kl+1},u(\boldsymbol{X}_{kl+1}),Child(\boldsymbol{X}_{kl+1}),u\bigl{(}Child(\boldsymbol{X}_{kl+1})\bigr{)}\Bigr{)}}}$ (16)
We now prove Lemma 3.2 by induction. By Definition 3.1, the globally optimal objective function cost is equal to the conditionally optimal cost of the last layer $G^{\star}(\boldsymbol{X}|U^{\star})=G^{*}(\boldsymbol{X}_{l_{\text{max}}}|U^{*}_{l_{\text{max}}},U^{*}_{<l_{\text{max}}})$. Note that conditional optimality of $U^{*}_{<l_{\text{max}}}$ is defined for a given cluster mapping at the specified layer, hence $U^{*}_{l_{\text{max}}}$ refers to the cluster mapping at layer $l_{\text{max}}$ that minimises $G$ globally. By equation (16), a cluster mapping with conditional optimality $G^{*}(\boldsymbol{X}_{l+1}|U^{*}_{l+1},U^{*}_{<l+1})$ at layer $l+1$ requires a conditionally optimal cluster mapping at layer $l$ Therefore, by induction, all layers of an optimal cluster mapping must themselves be conditionally optimal.
3.2.2 Branching
@@ 3.2.3 Optimality
DCMAP will incrementally explore branches (i.e. cluster mapping solutions) from layer 00 through $l_{\text{max}}$ to find every least cost $G^{\star}(\boldsymbol{X},U^{\star}_{i})$ solution.
@@
Each iteration of DCMAP corresponds to an exploration of a cluster-layer mapping $kl$. When an iteration produces a complete cluster mapping $\text{U}[b^{\prime},\boldsymbol{X}_{l}]$ for layer $l$ that is the same as one or more branches $b^{\prime\prime}$, a link is made between these branches in the same iteration. By Definition 3.1 of conditional optimality, any dominated branches sharing the same cluster mapping for that layer $\text{U}[b^{\prime},\boldsymbol{X}_{l}]$ can be pruned since by Lemma 3.2, an optimal cluster mapping comprises layer-by-layer conditionally optimal cluster mappings. Pruning of all dominated linked branches happens at the beginning of an iteration (Component Prune Linked Dominated Branches), so no iterations are wasted exploring a dominated branch.
@@
Therefore, it is not possible to find, using an equally informed algorithm, all the least cost cluster mappings in fewer iterations than DCMAP in the absence of branch activation. Examples of alternate algorithms for cluster mapping include tree search algorithms like breadth first or depth first search [16]. However, the number of iterations needed to find the first optimal solution is also important.
Consider the impact of pruning on search iterations. By Definition 3.1 of conditional optimality, pruning of dominated branches can only occur when a layer is completed (Component Layer Updates). Therefore, changing the order of updating of cluster-layers through branch activation in DCMAP has no effect on the total number of search iterations; it is already optimal (Lemma 3.4). However, branch activation, which evaluates one search branch to completion at a time, can potentially reduce the number of iterations to find the first least cost solution. This occurs as cluster-layers in a branch are updated in a semi-greedy ordering since each iteration evaluates cluster-layer proposals with the lowest $\hat{g}(\boldsymbol{X}_{i})$ first (Component Compute Costs).
@@
$\overline{N}_{T}=\overline{N}\overline{N}_{K}\overline{N}_{P}.$ (17)
@@
4 Case Study
We use as a case study the complex systems DBN model for seagrass resilience, described briefly in Section 1. This model, which has 25 discrete nodes per time-slice (Fig. 3) and 96 time slices [26], captures biological, ecological and environmental dynamics and their interactions to predict the impact on resilience of a series of dredging stressors and its cumulative effects over time. It demonstrates small world or power law connectivity (Table 3 and 4) with a few nodes having comparatively many parent or child nodes, and the majority having few [25]; most nodes had two to four states, with a few having ten or more states. [26] incorporated expert knowledge and multiple data sources (e.g. experimental and monitoring data) to inform the model; however, the computation time prohibited the use of Bayesian inference, which enables better data integration and estimation of uncertainty.
Table 3: Distribution of in- and out-degree (degree and number of nodes with that degree) in the one time-slice DBN.

### v6

@@ 3. Algorithm and Analysis
3.1. Algorithm Components
Our method uses $\hat{g}(\boldsymbol{X}_{kl})$ to prioritise cluster-layer proposals to update (i.e. pop), and $G(\boldsymbol{X}_{l})$ to facilitate local optimisation by identifying and cutting off further exploration of dominated solutions. It iterates from layer 0 (i.e. leaf nodes) through to $l_{\text{max}}$, storing unique and non-dominated cluster mappings on separate search branches. The algorithm terminates if, given current cluster-layer proposals on the queue and additive costs, the current costs of these potential solutions exceed the least-cost solutions found so far. Potentially, near-optimal solutions can be returned at earlier iterations of the search.
3.1.1. Data Structures and Initialisations
@@
[…]
@@
$\begin{aligned} \text{Br}_{\text{lk}}\in\\ \mathbb{R}^{B\times l_{\text{max}}\times B\times l_{\text{max}}}\end{aligned}$ $\text{Br}_{\text{lk}}[i,l,j,l]=1$ indicates a link at layer $l$ between branches $i$ and $j$.
@@
$Q$, $Q^{\prime}$ A queue with entries $q=\{b,u(\boldsymbol{X}_{kl}),l,\hat{g}(\boldsymbol{X}_{kl}),G(\boldsymbol{X}_{l})\}$. We denote $Q^{\prime}$ as eligible entries for updating (popping) $q\in Q:\text{Br}_{\text{pl}}[b]\geq q\mathdollar l$, aka activated branches.
@@
5:end procedure
3.1.2. Pop Queue or Activate Branch
Potentially, $Q$ could be treated as a heap to maximise computational efficiency (Lavalle2006). We use $Pop()$ to refer to removing the lowest cost queue entry, choosing between equicost entries arbitrarily (Component Pop Queue or Activate Branch). If there are no eligible entries on the queue, i.e. no branch has $\text{Br}_{\text{pl}}\geq l$, then a branch is chosen for activation. This makes queue entries on that branch potentially eligible. Activation randomly switches between using $Peek()$ to return, but not remove, the lowest cost entry, or $Peek^{\prime}()$ to return an entry at the lowest layer on the queue, choosing arbitrarily between entries at the same layer. This facilitates switching between best-first and breadth-first search (Theorem 3.4). After activation, DCMAP tries again to pop the queue if it is not empty; the algorithm only proceeds to the next step (Component Prune Linked Dominated Branches) if a queue entry was successfully popped.
@@ 3.1.2. Pop Queue or Activate Branch
5:   $\text{Br}_{pl}(q\mathdollar b)\leftarrow|\text{Br}_{pl}(q\mathdollar b)|$
6:  end if
7:end procedure
3.1.3. Prune Linked Dominated Branches
@@
4:end procedure
3.1.4. Propose Cluster Mappings
The requirement for contiguous clusters (Definition 3) means that a node can only be mapped to a new cluster $u_{\text{new}}$, or the cluster of one or more child nodes $\bf{u}_{\text{exist}}$. For the former, we can ensure uniqueness and avoid accidentally mapping to an existing cluster by associating a unique label $1..n$ with each node. For the latter, define a priori a $n\times n$ matrix $S:S[i,j]=\{0,1\}$, where $1$ indicates that node $X_{i}$ can potentially be mapped to the same cluster as node $X_{j}$. To ensure contiguity, Algorithm 2 works backward by identifying, for a given node $X_{i}$, which nodes $X_{j}$ at layer $l(X_{i})$ or earlier layers can be in the same cluster as $X_{i}$ that are connected by a common parent node. Note that descendants are child nodes, child of child nodes and so on. For node $X_{j}$, $u(X_{i^{\prime}}):S[i^{\prime},j]=1$ are all the node clusters that $X_{j}$ could be mapped to. To preserve contiguity, node $X_{j}$ must be such that there are no paths $\boldsymbol{X}_{p}$ following the directed arcs between $X_{i}$ and $X_{j}$ containing any other nodes i.e. $\boldsymbol{X}_{p,ij}=\{X_{i},X_{j}\}$ or $\boldsymbol{X}_{p,ij}=\varnothing$. The former implies $X_{j}$ is a parent of $X_{i}$, and the latter is where both nodes are at the same layer, $l(X_{j})=l(X_{i})$.
@@ 3.1.4. Propose Cluster Mappings
When applied iteratively and incrementally from layer 0 to $l_{\text{max}}$ (Fig. 2), all feasible contiguous cluster mappings (Definition 3) can be generated using matrix $S$ (Algorithm 2) to identify all nodes $X_{j}$ that can be mapped to the same cluster as node $X_{i}$.
@@
The assumption of iterating incrementally over $l=0,1,\ldots,l_{\text{max}}$ means that at layer $l$, only nodes at layer $l$ or earlier need to be considered for mapping to the same cluster. Additionally, Algorithm 2 only allows nodes $X_{i}$ and $X_{j}$ to be mapped to the same cluster if they are directly connected (i.e. a parent or child of the parent) with no paths going through an intermediary node; this guarantees contiguity, even if $X_{i}$ and $X_{j}$ are not in successive layers (Remark 2.1). At each iteration, feasible cluster mappings for node $X_{j}$ are the clusters of nodes $X_{i}$ associated with non-zero rows of $S[,j]$, which might be in the same layer or at a child layer. By Definition 1 of layers, there must exist at least one parent-child path between each node in one layer and the previous layer, hence, there are no contiguous cluster mappings that cannot be generated by iteratively and incrementally applying $S$ from $l=0$ to $l=l_{\text{max}}$. ∎
@@
(13) $\footnotesize\begin{bmatrix}&A&B&C&D&E&F&G\\ A&0&0&0&0&0&0&0\\ B&0&0&0&0&0&0&0\\ C&0&0&0&0&0&0&0\\ D&1&1&0&0&0&0&0\\ E&0&0&1&0&0&0&0\\ F&1&0&0&0&0&0&1\\ G&0&0&0&1&1&0&0\\ \end{bmatrix}$
For instance, node $F$ (row $F$) could share the same cluster as nodes $A$, $D$ and itself (Algorithm 2). Thus, when updating (popping) node $D$ (column $D$) as the search progresses from layer 0 to $2$, it could be mapped to the current cluster of node $F$ or $G$ or itself. Locally, this could lead to contiguous clusters involving D: $(A,F,D,G)$, $(A,F,D)$, $(F,D)$, $(D,G)$, $(D)$.
The above constrained space of feasible cluster mappings could be pre-computed at initialisation and are enacted in Component Update Possible Cluster Mappings and Queue through $\hat{\textrm{U}}$. Component Propose Cluster Mappings generates from $\hat{\textrm{U}}$ and particular $k$ and $l$ the cluster-mappings for cost evaluation at each iteration. Intuitively, for a given branch and layer $l$, $\boldsymbol{Z}_{1}$ represents the nodes that cannot be in any other cluster $k^{\prime}\neq k$, $\boldsymbol{Z}_{3}$ are nodes that have already been mapped to another cluster $k^{\prime}$, and $\boldsymbol{Z}_{2}$ are the nodes that could be in $k$ or another cluster $k^{\prime}$. Each of the $N$ unique cluster mapping combinations are stored on separate branches, producing N1N-1 new branches, done by updating data structures and duplicating queue entries from the current branch $b$ for new branches $b^{\prime}$.
@@
6:   Duplicate $q^{\prime}\in Q:q^{\prime}\mathdollar b=b,q^{\prime}\mathdollar l=l$ with $q^{\prime}\mathdollar b=b^{\prime}$
7:  end for
8:end procedure
9:procedure generatecombos($\boldsymbol{Z},\text{U},\hat{\text{U}}$)
10:  $\boldsymbol{Z}_{2}\leftarrow\boldsymbol{Z}\setminus\{\boldsymbol{Z}_{1},\boldsymbol{Z}_{3}\}:\bigl(\text{U}[b,\boldsymbol{Z}_{3}]\neq u(\boldsymbol{X}_{kl}),U[b,\boldsymbol{Z}_{3}]\neq 0\bigr),\bigl(\forall\hat{\text{U}}[b,u\neq u(\boldsymbol{X}_{kl}),\boldsymbol{Z}_{1}]=0\bigr)$
11:  $\boldsymbol{X}_{i=1..N}\leftarrow\Bigl\{\boldsymbol{z}_{j}\cup\boldsymbol{Z}_{1}\text{ }\forall\boldsymbol{z}_{j}\in\text{Combinations}{\boldsymbol{Z}_{2}\choose 1..|\boldsymbol{Z}_{2}|},\boldsymbol{Z}_{1}\Bigr\}$
12:end procedure
@@
(14) $\hat{\textrm{U}}\left[b,l(\boldsymbol{X}_{i})=l\right]=\begin{bmatrix}k&X_{1}&X_{2}&X_{3}\\ 2&0&1&1\\ 3&1&1&1\\ \end{bmatrix}$
@@ 3.1.6. Layer Updates
6:  end if
7:end procedure
3.1.7. Update Possible Cluster Mappings and Queue
@@
5:end procedure
3.2. Theoretical Analysis
This section proves analytically that DCMAP is optimal in the sense that it finds all least cost cluster mappings, and does so such that an equally informed algorithm cannot do so in fewer iterations. This is achieved through conditional optimality (derived in Section 3.2.1), the exploration of all possible contiguous cluster mappings through branching (Section 3.2.2), and combined to show the optimality of DCMAP through evaluating non-dominated branches, i.e. potential cluster mapping solutions, at the earliest opportunity (Section 3.2.3). An additional property is derived for super additivity of cost functions for a heuristic to further speed up the search (Appendix B).
3.2.1. Conditional Optimality
@@
Following Definition 6, we say that $U^{*}_{<l}$ dominates $U^{i}_{<l}$ if $G^{*}(\boldsymbol{X}_{l}|U_{l},U^{*}_{<l})<G(\boldsymbol{X}_{l}|U_{l},U^{i}_{<l})$.
@@
(16) ${\sum_{k=1}^{m_{l+1}}{c\Bigl(\boldsymbol{X}_{kl+1},u(\boldsymbol{X}_{kl+1}),Child(\boldsymbol{X}_{kl+1}),u\bigl(Child(\boldsymbol{X}_{kl+1})\bigr)\Bigr)}}$
We now prove Lemma 3.2 by induction. By Definition 7, the globally optimal objective function cost is equal to the conditionally optimal cost of the last layer $G^{\star}(\boldsymbol{X}|U^{\star})=G^{*}(\boldsymbol{X}_{l_{\text{max}}}|U^{*}_{l_{\text{max}}},U^{*}_{<l_{\text{max}}})$. Note that conditional optimality of $U^{*}_{<l_{\text{max}}}$ is defined for a given cluster mapping at the specified layer, hence $U^{*}_{l_{\text{max}}}$ refers to the cluster mapping at layer $l_{\text{max}}$ that minimises $G$ globally. By equation (16), a cluster mapping with conditional optimality $G^{*}(\boldsymbol{X}_{l+1}|U^{*}_{l+1},U^{*}_{<l+1})$ at layer $l+1$ requires a conditionally optimal cluster mapping at layer $l$ . Therefore, by induction, all layers of an optimal cluster mapping must themselves be conditionally optimal.
3.2.2. Branching
@@ 3.2.3. Optimality
DCMAP will incrementally explore branches (i.e. cluster mapping solutions) from layer 0 through $l_{\text{max}}$ to find every least cost $G^{\star}(\boldsymbol{X},U^{\star}_{i})$ solution.
@@
Each iteration of DCMAP corresponds to an exploration of a cluster-layer mapping $kl$. When an iteration produces a complete cluster mapping $\text{U}[b^{\prime},\boldsymbol{X}_{l}]$ for layer $l$ that is the same as one or more branches $b^{\prime\prime}$, a link is made between these branches in the same iteration. By Definition 7 of conditional optimality, any dominated branches sharing the same cluster mapping for that layer $\text{U}[b^{\prime},\boldsymbol{X}_{l}]$ can be pruned since by Lemma 3.2, an optimal cluster mapping comprises layer-by-layer conditionally optimal cluster mappings. Pruning of all dominated linked branches happens at the beginning of an iteration (Component Prune Linked Dominated Branches), so no iterations are wasted exploring a dominated branch.
@@
Therefore, it is not possible to find, using an equally informed algorithm, all the least cost cluster mappings in fewer iterations than DCMAP in the absence of branch activation. Examples of alternate algorithms for cluster mapping include tree search algorithms like breadth first or depth first search (Lavalle2006). However, the number of iterations needed to find the first optimal solution is also important.
Consider the impact of pruning on search iterations. By Definition 7 of conditional optimality, pruning of dominated branches can only occur when a layer is completed (Component Layer Updates). Therefore, changing the order of updating of cluster-layers through branch activation in DCMAP has no effect on the total number of search iterations; it is already optimal (Lemma 3.5). However, branch activation, which evaluates one search branch to completion at a time, can potentially reduce the number of iterations to find the first least cost solution. This occurs as cluster-layers in a branch are updated in a semi-greedy ordering since each iteration evaluates cluster-layer proposals with the lowest $\hat{g}(\boldsymbol{X}_{i})$ first (Component Compute Costs).
@@
(17) $\overline{N}_{T}=\overline{N}\overline{N}_{K}\overline{N}_{P}.$
@@
4. Case Study
We use as a case study the complex systems DBN model for seagrass resilience, described briefly in Section 1. This model, which has 25 discrete nodes per time-slice (Fig. 3) and 96 time slices (Wu2018), captures biological, ecological and environmental dynamics and their interactions to predict the impact on resilience of a series of dredging stressors and its cumulative effects over time. It demonstrates small world or power law connectivity (Table 3 and 4) with a few nodes having comparatively many parent or child nodes, and the majority having few (Watts1998); most nodes had two to four states, with a few having ten or more states. (Wu2018) incorporated expert knowledge and multiple data sources (e.g. experimental and monitoring data) to inform the model; however, the computation time prohibited the use of Bayesian inference, which enables better data integration and estimation of uncertainty.
Table 3. Distribution of in- and out-degree (degree and number of nodes with that degree) in the one time-slice DBN.